40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Entry  Mon Apr 22 09:54:21 2019, Joe, General, , Shortening cavity (A5,A14,PZT11,PZT22) to get closer to design FSR CM1_IMG_7699.jpgCM2_IMG_7697.jpg
    Reply  Mon Apr 22 19:54:28 2019, Koji, General, , OMC(004): Spot positions at the end of Apr 22nd misalignment4.pdf
Message ID: 351     Entry time: Mon Apr 22 09:54:21 2019     Reply to this: 352
Author: Joe 
Type: General 
Category:  
Subject: Shortening cavity (A5,A14,PZT11,PZT22) to get closer to design FSR 

[Koji,Joe,Philip,stephen]

in units 20um per div on the micrometer [n.b. we reailised that its 10um per div on the micrometer]

CM1 inner screw pos: 11.5

cm1 outer screw pos: 33.5

cm2 inner screw pos: 11

cm2 outer screw pos: 13

the cavity is currently 3mm too long, move each mirror closer by 0.75mm

CM1 inner screw pos: 11.5+37.5 = 49

cm1 outer screw pos: 33.5+37.5= 71

cm2 inner screw pos: 11+37.5 = 48.5

cm2 outer screw pos: 13+37.5 = 50.5

The screws on the micrometers were adjusted to these values.

cleaned cm1 (PZT 11). There was a mark near the edge which we were not able to remove with acetone. On the breadboard there were 3 spots which we could not remove with acetone. Once we wiped the mirror and breadboard we put the mirror back. 

FM2 (A5). The prism looked quite bad when inspected under the green torch, with lots of lines going breadthways. We thought about replacing this with A1, however this has had the most exposure to the environment according to koji. This has a bit of negative pitch, so would bring down the beam slightly. We decided to continue to use A5 as it had worked fairly well before. The breadboard was cleaned, we could see a few spots on it, they were cleaned using acetone.

FM1 (A14). Near the edge of the bottom surface of the prism we could see some shiny marks, which may have been first contact. We attempted to scrape them off we tweezers. The breadboard looked like it had a few marks on it. These were hard to remove with the acetone, it kept leaving residue marks. We used isopropanol to clean this now, which worked much better. The sharp edges of the breadboard can cause the lens tissue to tear a bit, so it took a few rounds of cleaning before it looked good to put a prism on. The mirror was put back onto the breadboard. 

The cavity was aligned, then we realised that 1 turn is 500um, so its still too long (1.75mm long). The FSR was 264.433Mhz, which is 

CM2 still showed quite a bit more scattering than CM1, so we want to move this beam.

CM1:

  • inner = 0.405mm
  • outer = 0.67mm

CM2

  • inner = 0.507mm
  • outer = 0.42mm

want to increase by 1.7/4 = 0.425, so

CM1:

  • inner = 0.405+ 0.425 mm = 0.83 mm
  • outer = 0.67+ 0.425mm = 1.095 mm

CM2

  • inner = 0.507 + 0.425mm = 0.932 mm
  • outer = 0.42 + 0.425mm = 0.845 mm

we tried to align the cavity, however the periscope screws ran out of range, so we changed the mircometers on CM2. We tried this for quite some time, but had problems with the beam reflected from the cavity clipping the steering mirror on the breadboard (to close to the outer edge of the mirror). This was fixed by changing the angle of the two curved mirrors. (We should include a diagram to explain this).

The cavity was locke, the FSR was measured using the detuned locking method, and we found that the FSR = 264.805 MHz, which corresponds to a cavity length of 1.1321m

we took some photos, the spot is quite far to the edge of the mirrors (3 to 4mm), but its near the centre vertically. photos are 

123-7699 = CM2

123-7697 = CM1

 

 

Attachment 1: CM1_IMG_7699.jpg  2.473 MB  Uploaded Mon Apr 22 20:53:49 2019  | Show | Hide all | Show all
Attachment 2: CM2_IMG_7697.jpg  3.257 MB  Uploaded Mon Apr 22 20:53:55 2019  | Hide | Hide all | Show all
CM2_IMG_7697.jpg
ELOG V3.1.3-